The length of a rectangle is 3 m more than twice its width, IF the perimeter of the rectangle is 42 m, find the dimensions of the rectangle.

Answer by checkley71(8403)

L=2W+3
42=2(2W+3)+2W
42=4W+6+2W
42=6W+6
6W=42-6
6W=36
W=36/6
W=6M FOR THE WIDTH.
L=2*6+3
L=12+3
L=15M FOR OF THE LENGTH.
PROOF
42=2*15+2*6
42=30+12
42=42